OpenStudy (anonymous):

Help me find the solution to x^3-x^2-2x=12?

OpenStudy (anonymous):

Find the solutions to this equation algebraically using the Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem.

OpenStudy (anonymous):

http://www.easycalculation.com/algebra/cubic-equation.php x=3

OpenStudy (anonymous):

i need it in each equation i thinknd it has to shdow work

OpenStudy (anonymous):

Hint: \[x^3-x^2-2 x-12=(x-3) \left(x^2+2 x+4\right) \]

OpenStudy (anonymous):

can u show ur work ?

OpenStudy (anonymous):

Do you know remainder theorum? Here is a great site that also shows you a quick shortcut for the division http://www.purplemath.com/modules/remaindr.htm

OpenStudy (anonymous):

Used Mathematica that may have used the factor theorem http://www.purplemath.com/modules/factrthm.htm

OpenStudy (anonymous):

huh? how?

OpenStudy (anonymous):

called synthetic division. You have to guess the value for x and then do the division, if you end up with no remainder then thats a factor use that factor to get to the satge robtobey has with a factor and a quadratic. Now factor the quadratic to get the final answer in the form of (x+?)(x+?)(x+?)

OpenStudy (anonymous):

start with one and keep going one at a time you will fond 3 is a factor

OpenStudy (anonymous):

i have to use the things above or i dont get credit.....

OpenStudy (nali):

factor then when u solve for x, u will probabily use teh rest of the things u listed above

OpenStudy (anonymous):

yep, looking them up you will use most of them just by using the remainder theorum and the division of polynomials, thus proving that the number of solution are the same as the degree of the poly. test your answer usingn the synthetic divison as it makes you life easy.

OpenStudy (anonymous):

step 1 is definitely to find a factor of the poly x^3